# Distributed Saddle-Point Dynamics in Multilayer Networks

### v1

@@ I Introduction
The quantitative study of networks has provided deep understanding of the structure and dynamics of complex systems in diverse fields [1, 2], such as biology [3], neuroscience [4], engineering [5, 6], and social sciences [7]. However, the traditional single-layer network formalism does not capture the full complexity of real-world systems where multiple types of interaction occur simultaneously across distinct layers [8]. Multilayer networks provide a more holistic framework for representing these interconnected systems, allowing researchers to explore a broader range of dynamics and interactions that are not apparent in single-layer or multiplex network models [9].
@@
Given the multidimensional nature of these processes, it is worth establishing a tensorial representation for such systems. Tensors generalize scalars, vectors, and matrices [20] and provide a concise framework for modeling and solving multidimensional problems. Tensors are widely used in all fields, from linear algebra and physics to engineering [21, 22]. In  [23] it is developed a tensor-based mathematical formalism that provides a general, robust, and comprehensive framework for representing multilayer networks with all their complexities, offering insights that remain hidden when using traditional adjacency matrices. This formalism has been applied explicitly to study disease localization [24], generalization of the concept of assortativity [25], spreading [26], and eigenvector multicentrality [27].
@@
Despite the recent advances in multilayer networks, progress in generalizing distributed optimization problems from single-layer to multilayer networks has been limited. In conventional distributed optimization, each node typically has access to its own convex function, and the overall objective is defined as the sum of these individual functions [28]. To solve this, each agent seeks the optimal solution to the global convex function by exchanging information with neighboring nodes and performing local computations [29]. A variety of methods have been developed to address this problem, ranging from discrete-time gradient descent approaches [29] to more recent continuous-time techniques [30, 31, 32]. In particular, from a continuous-time perspective, optimization algorithms have been framed as dynamical systems [33], with methods based on saddle-point dynamics [34, 35] gaining prominence.
@@
In a previous work [36, 37], we developed distributed optimization algorithms tailored to multiplex networks, leveraging the supra-Laplacian matrix to handle the diffusion dynamics within and across layers. Although these algorithms demonstrated success in modeling systems with distinct but fixed interlayer connections, such as in infrastructure networks, they are limited in their applicability to more complex multilayer systems. Specifically, multiplex networks assume identical sets of nodes across layers and restrict interlayer connections to counterpart nodes, a limitation when modeling systems with heterogeneous interlayer interactions. To overcome these limitations, a more generalized framework is necessary to handle complex interactions in multilayer networks. The tensorial formalism introduced in [23] offers a powerful approach to capture both intra- and interlayer dynamics within these systems. By generalizing the network descriptors and dynamics to tensorial representations, this framework enables us to model a richer class of multilayer networks where interactions across layers are not restricted to node counterparts, but can vary in structure and strength.
@@
In this paper, we extend our previous work on distributed optimization in multiplex networks by incorporating the tensorial framework for multilayer networks. This generalization allows us to handle a broader class of systems with complex interlayer connections. We propose novel distributed optimization algorithms based on tensorial dynamics, analyze their convergence properties, and validate them through numerical examples.
@@
The main contributions of this work are as follows: (1) extending the consensus and diffusion equations to a tensor-based representation for multilayer networks, thereby generalizing the relationship between diffusion processes in multiplex and multilayer networks and consensus-driven distributed optimization algorithms; (2) formulating a distributed saddle-point algorithm tailored for convex optimization in general multilayer networks; and (3) proving the convergence of each node in every layer to the optimal value of the convex objective function. In addition, we examine the role of interlayer diffusion and its effect on optimization efficiency in multilayer networks, illustrating the applicability of our framework in real-world scenarios such as multi-energy systems. We further examine the occurrence of critical phenomena in consensus time by varying the diffusion constants. In addition, several numerical examples are presented to validate the proposed algorithms and investigate the influence of interlayer diffusion on the consensus time.
@@ II-A2 Multilayer Networks
are the fourth-order (i.e. rank-4) tensors of the canonical basis in the space $\mathbb{R}^{N\times N\times M\times M}$. We have adopted the notation $N\equiv N(M)$ to remark that the number of nodes in each layer might be different. The multilayer adjacency tensor $M^{\alpha\tilde{\mu}}_{\beta\tilde{\nu}}$ is a versatile construct that can represent a wide range of complex relationships between nodes across different layers.
@@ II-B Diffusion processes on multilayer networks
II-B1 Diffusion on single-layer networks
Using the tensor notation, the diffusion processes on single-layer networks can be described as follows. Let $x_{\alpha}(t)$ be a state vector of nodes at time $t$. The diffusion equation is given by
@@ II-B1 Diffusion on single-layer networks
II-B2 Diffusion on multilayer networks
In this case, a diffusion equation for multilayer networks must incorporate terms that account for interlayer diffusion. Let $X_{\alpha\tilde{\mu}}(t)$ denote the state tensor of nodes in each layer at time $t$. Thus, a diffusion equation for a multilayer network is obtained as follows,
@@ II-B2 Diffusion on multilayer networks
Similar to the single-layer case, it is presented a multilayer combinatorial Laplacian tensor as
@@ III Problem Statement
$f_{\alpha\tilde{\mu}}(x)=\sum_{\tilde{h}=1}^{M}\sum_{i=1}^{N(M)}f_{i\tilde{h}}(x)e_{\alpha}(i)e_{\tilde{\mu}}(\tilde{h}),$ (16)
@@ VI Conclusion
In this paper, we develop a distributed optimization algorithm tailored for generalized multilayer networks within a tensorial framework. By extending the connection between diffusion and consensus dynamics and leveraging control approaches used in multiplex networks, we propose a distributed tensorial primal-dual saddle-point algorithm that achieves fast convergence and robustness. Similar to multiplex networks, we find that intra- and interlayer diffusion constants serve as control parameters for optimal consensus time. Each layer first reaches a local consensus before achieving global consensus across layers.
@@
Possible extensions of this work involve obtaining a diffusion operator for more general higher-order networks and the application of the multilayer optimization proposed approach [40]. A natural continuation of this work is to study hypergraphs consensus dynamics [41] and consensus in simplicial complexes [42, 43]. On the other hand, topological Dirac operators offer a generalization scenario which could be exploited to reveal new properties in complex systems, including new consensus dynamics [44]. Another potential extension involves the introduction of disorder [45, 46] and the application of graph zeta functions to analyze diffusion dynamics in higher-order networks [47].
@@ Appendix B Proof of Theorem 1
First, consider the convergence analysis to the tensor points $(Y^{\star}_{\beta\tilde{\nu}},\Lambda^{\star}_{\beta\tilde{\nu}})$. Let $\tilde{Y}_{\beta\tilde{\nu}}=Y_{\beta\tilde{\nu}}-Y^{\star}_{\beta\tilde{\nu}}$ and $\tilde{\Lambda}_{\beta\tilde{\nu}}=\Lambda_{\beta\tilde{\nu}}-\Lambda^{\star}_{\beta\tilde{\nu}}$. In the new variables, the saddle-point flow (20)-(21) yields
@@
To demonstrate the global asymptotic stability of the equilibrium tensor point $(Y^{\star},\Lambda^{\star})$, we apply LaSalle’s invariance principle, showing that there are no trajectories in $E$ other than the equilibrium point $\tilde{Y}=0$ and $\tilde{\Lambda}=0$. It is observed that if $\tilde{Y}=0$ the system (30)-(31) casts to
@@
By examining the dual dynamics (31), we can observe that $U^{\alpha\tilde{\mu}}\dot{\Lambda}_{\alpha\tilde{\mu}}=0$. Therefore, we obtain $\text{average}(\Lambda(t))=\text{average}(\Lambda_{0})$ $\forall t\geq 0$. As a consequence, the convergence to a single tensor saddle-point $(Y^{\star},\Lambda^{\star})$ of the flow (30)-(31) is guaranteed, which satisfies $Y^{\star}_{\alpha\tilde{\mu}}=x^{\star}U_{\alpha\tilde{\mu}}$ and $\Lambda^{\star}_{\alpha\tilde{\mu}}=\bar{\Lambda}_{\alpha\tilde{\mu}}+\text{average}(\Lambda_{0})U_{\alpha\tilde{\mu}}$. ∎

### v2

@@ I Introduction
The quantitative study of networks has provided a deep understanding of the structure and dynamics of complex systems in various fields [1, 2], such as biology [3], neuroscience [4], engineering [5, 6], and social sciences [7]. However, the traditional single-layer network formalism does not capture the full complexity of real-world systems where multiple types of interaction occur simultaneously across distinct layers [8]. Multilayer networks provide a more holistic framework for representing these interconnected systems, allowing researchers to explore a broader range of dynamics and interactions that are not apparent in single-layer or multiplex network models [9].
@@
Given the multidimensional nature of these processes, it is worth establishing a tensorial representation for such systems. Tensors generalize scalars, vectors, and matrices [20] and provide a concise framework for modeling and solving multidimensional problems. Tensors are widely used in all fields, from linear algebra and physics to engineering [21, 22]. In  [23] it is developed a tensor-based mathematical formalism that provides a general, robust, and comprehensive framework to represent multilayer networks with all their complexities, offering insights that remain hidden when using traditional adjacency matrices. This formalism has been applied explicitly to study disease localization [24], generalization of the concept of assortativity [25], spreading [26], and eigenvector multicentrality [27].
@@
Despite recent advances in multilayer networks, progress in generalizing distributed optimization problems from single-layer to multilayer networks has been limited. In conventional distributed optimization, each node typically has access to its own convex function, and the overall objective is defined as the sum of these individual functions [28]. To solve this, each agent seeks the optimal solution to the global convex function by exchanging information with neighboring nodes and performing local computations [29]. Various methods have been developed to address this problem, ranging from discrete-time gradient descent approaches [29] to more recent continuous-time techniques [30, 31, 32]. In particular, from a continuous-time perspective, optimization algorithms have been framed as dynamical systems [33], with methods based on saddle-point dynamics [34, 35] gaining prominence.
I-A Related Work
A multilayer network is formed by several interacting networks [36], [9]. The first introduction of multilayer networks is in the context of social science to characterize the different types of social relationships that exist between the nodes of a social network [37]. Currently, multilayer networks are investigated in many fields, including diffusion dynamics and information spreading [38], [39], [40], epidemic spreading [41], neuroscience [42], molecular biology [43], ecology [10], and economy [44]. Multilayer networks have also been applied effectively across various engineering domains to model and analyze complex systems with multiple interacting components such as power networks, where the interdependence between power grids and communication infrastructures is effectively captured using multilayer network models [16]. One layer represents the electrical grid, while another represents the communication network that monitors and controls the grid’s operations. This modeling enables the analysis of cascading failures [50], [51], where a disruption in the power grid can impact the communication network and vice versa, aiding in the design of more resilient interdependent infrastructures. Another successful application is urban transportation networks including roads, railways, buses, and metro systems, which can be represented as multilayer networks where each mode of transport forms a layer [45]. Information propagation is also a relevant application of multilayer networks [46]. Although there are significant advances in other engineering applications phenomena such as percolation [47], there is a lack of developments in the generalization of single-layer network distributed optimization problems to multilayer networks.
A specific type of multilayer network is multiplex networks, where there is a one-to-one correspondence between each node across all layers. In other words, the same set of nodes is connected by different types of interactions in each layer. A multiplex network is a set of graphs composed by different layers in which the same set of nodes may be connected to each other by means of links belonging to different classes or types [48]. Multiplex networks are initially introduced in the context of social networks, where individuals are linked through various social relationships such as friendship, collaboration, and family, or communicate through different channels such as email, mobile phones, and instant messaging [49]. Over time, the concept of multiplex networks has been extended to encompass a wide range of domains, including transportation systems [50], and brain networks [51].
Multilayer networks represent a generalization of peer-to-peer architectures, where agents are embedded within single-layer communication networks. In peer-to-peer architectures, consensus protocols have garnered significant attention and research interest due to their applications in synchronization, formation control, learning, and the coordination of multi-robot tasks, among others. These systems are inherently distributed.
Distributed computational models and coordination mechanisms are based on local communication between agents. Such models are crucial for advancing technologies like automated transportation, smart cities, smart power grids, and robotic systems. At the heart of these technologies are devices capable of collecting data (e.g., gathering measurements from their environment), processing information, and communicating with one another to share estimates, thereby ensuring optimal and reliable system performance. Learning methods play a fundamental role in data processing, while communication between devices enables collaborative learning from the collective data within the system, without the need to share the data stored locally on the devices [52].
A common framework for addressing these challenges is provided by distributed computational models, particularly those incorporating machine learning, which utilize consensus mechanisms as virtual coordinators. These mechanisms facilitate coordination and information exchange without centralized control, making them essential for scalable and efficient distributed system operation. One of the most robust tools to generate consensus mechanisms to coordinate agents are the distributed consensus-based gradient methods for convex learning problems over static graphs. The goal of these methods is to collaboratively solve the following problem:
$\min_{x\in\mathbb{R}^{p}}f(x):=\sum_{i=1}^{n}f_{i}(x)$ (1)
where $x$ is the global decision variable and each function $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ is convex and known by agent $i$ only. The agents are embedded in a communication network, and their goal is to obtain an optimal and consensual solution through local neighbor communications and information exchange. This local exchange is desirable in situations where the exchange of a large amount of data is prohibitively expensive due to limited communication resources. Since the problem (1) is over a networked system of agents $n$, the way to characterize this networks is through the definition of a graph $G$ as a tuple $G=([m],\mathcal{E})$ of a set of agents $[n]=\{1,2,3,...,n\}$ and a set of edges =$\mathcal{E}=\{(i,j)\quad\forall\,i,j\in[n]\}$. We write $(i,j)$ to represent an edge connecting agents $i$ and $j$. The agents want to solve the problem (1) collaboratively in the sense that they are willing to share some estimates with their immediate neighbors in the graph, but they are not willing (or not allowed) to share their data, which essentially means they do not reveal their functions $f_{i}$ [52].
Given the communication graph $G=([m],\mathcal{E})$, the problem is usually reformulated as follows:
$\min_{x\in\mathbb{R}^{p},i\in[n]}\sum_{i=1}^{n}f_{i}(x_{i})$
[…]
This reformulation is obtained from (1) by assigning a copy $x_{i}$ of the decision variable $x$ to each agent and imposing the requirement that all copies be the same. This reformulation serves as a departure point for discussing distributed methods that use consensus-type updates to enforce the equal-decision constraints in (2[52].
Since each $f_{i}$ depends on its own variable $x_{i}$, the objective function in (2) is decoupled, but these variables are coupled through edge-based constraints. The idea is to distribute the problem among the agents, by allowing each agent to know its neighbors in the graph. Using the local agent knowledge of the graph and functions, every agent $i$ can solve its own local part of the overall problem. However, for the agents to collectively solve the overall problem, each agent needs to align its variables with the variables $x_{j}$ corresponding to its neighbors $j\in\mathcal{N}_{i}$, being $\mathcal{N}_{i}=\{j:(i,j)\in\mathcal{E}\}$. As a mechanism for alignment of the variables, a consensus algorithm is used. The consensus algorithm is a distributed method that the agents can use to asymptotically agree on a decision vector. The distributed optimization algorithm updates by using a gradient of its private function, and is local in the sense that it relies on local information exchange. This method if often referred as consensus-based gradient method, due to its use of a mixing step that resembles the distributed consensus process. There are extensions and variations of this method for systems embedded in single-layer networks. For instance, in [53] it is proposed distributed gradient-based method (push–pull and G-push–pull) where each node maintains estimates of the optimal decision variable and the average gradient of the agents’ objective functions. In [54] the methods are extended to time-varying graphs. There are stochastic approaches to improve the methods [55, 56, 57, 58]. Furthermore, there are methods where the focus is in the gradient tracking, that is, each agent uses an auxiliary variable to asymptotically track the gradient of the global cost in expectation [59, 60, 61, 62, 63, 64, 65]. For the interested reader, a complete and concise tutorial can be found in [52] and references therein.
The natural extension of the discrete-time algorithms consists of a representation to continuous-time dynamics. To construct this representation, we start with the traditional consensus equation for single-layer networks of multiagent systems using the Laplacian matrix $L$. The Laplacian matrix also called the graph Laplacian is the conventional form to represent a graph and is defined as $L=D-A$, where $D$ is the degree matrix given by $d_{ij}=\text{deg}(n_{i})\delta_{ij}$, being $\text{deg}(n_{i})$ the degree of a node $n_{i}$ that counts the number of times an edge terminates at that node, and $A$ is the adjacency matrix given by $a_{ij}=1$ if $(i,j)\in\mathcal{E}$. The understanding of these matrices will be fundamental to the generalization to tensor which we shall explain in the next sections.
Now, the single-layer consensus-based dynamics can be presented in compact form as a diffusion equation
$\dot{x}=-Lx,$ (3)
where $L$ is the graph Laplacian. Using the consensus equation (3), several continuous-time saddle point dynamics have proposed to solve a distributed optimization problem in multiagent systems [34, 35, 66, 30, 67, 68, 33]
$\min_{x\in\mathbb{R}^{n}}\tilde{f}(x)=\sum_{i=1}^{n}f_{i}(x)\quad\text{s.t.}\quad Lx=0,$ (4)
where $f(x)$ is the global objective function to be minimized subject to the communication constraints between agents represented by the graph Laplacian. The corresponding saddle-point dynamics for optimization problem (4)
$\dot{x}$ $=$ $-Lx-\frac{\partial\tilde{f}(x)}{\partial x}-L\lambda,$ (5)
$\dot{\lambda}$ $=$ $Lx,$ (6)
where $\lambda$ is the vector of Lagrange multipliers [69]. The main idea is to take advantage of the continuous-time representations, and the diffusion dynamics in higher-order networked systems driven by Lapacian-like operators to generalize the dynamics given by (6), which are useful for single-layer networks, to multilayer networks.
I-B Summary of Contributions
@@
In a previous work [70, 71], we developed distributed optimization algorithms customized to multiplex networks, leveraging the supra-Laplacian matrix, $\mathcal{L}$, to handle diffusion dynamics within and between layers. Although these algorithms demonstrated success in modeling systems with distinct but fixed interlayer connections, such as in infrastructure networks, they are limited in their applicability to more complex multilayer systems. Specifically, multiplex networks assume identical sets of nodes across layers and restrict interlayer connections to counterpart nodes, a limitation when modeling systems with heterogeneous interlayer interactions. See Fig. 1.
Figure 1: Example of a multiplex network with two layers. Each layer has replica nodes, but a different topology. The connections between layers are one-by-one. On the side, we depict the Laplacians, $L_{1}$, $L_{2}$, the supra-Laplacian $\mathcal{L}$, and the optimal consensus dynamics for this system. The vector $y$ contains the state of each node within the two layers, while $\lambda$ is the vector of Lagrange multipliers. Further developments can be found in [70, 71].
To overcome these limitations, a more generalized framework is necessary to handle complex interactions in multilayer networks. The tensorial formalism introduced in [23] offers a powerful approach to capture intra- and interlayer dynamics within these systems. By generalizing the network descriptors and dynamics to tensorial representations, this framework enables us to model a richer class of multilayer networks where interactions across layers are not restricted to node counterparts, but can vary in structure and strength.
Figure 2: Example of a multilayer network with three layers. Each layer has different number of nodes and different topology. The connections between layers may not be one-by-one.
@@
In this paper, we extend our previous work on distributed optimization in multiplex networks by incorporating the tensorial framework for multilayer networks. This generalization allows us to handle a broader class of systems with complex interlayer connections (See Fig. 2). We propose novel distributed optimization algorithms based on tensorial dynamics, analyze their convergence properties, and validate them through numerical examples.
@@
The main contributions of this work are as follows: (1) extending the consensus and diffusion equations to a tensor-based representation for multilayer networks, thereby generalizing the relationship between diffusion processes in multiplex and multilayer networks and consensus-driven distributed optimization algorithms; (2) formulating a distributed saddle-point algorithm tailored for convex optimization in general multilayer networks; and (3) proving the convergence of each node in every layer to the optimal value of the convex objective function. In addition, we examine the role of interlayer diffusion and its effect on optimization efficiency in multilayer networks, illustrating the applicability of our framework in real-world scenarios such as multi-energy systems. We further examine the occurrence of critical phenomena in consensus time by varying the diffusion constants. In addition, several numerical examples are presented to validate the proposed algorithms and investigate the influence of interlayer diffusion on the consensus time. We can summarize the impacts and contributions of this work in Table I.
TABLE I: Comparison of proposed approach and related references
Ref. Network Type Optimization Method Limitations Our Contribution
[52, 67] Single-layer. Distributed Gradient descent: $\dot{x}=-Lx-\frac{\partial\tilde{f}(x)}{\partial x}-L\lambda,$ $\quad\dot{\lambda}=Lx$ Limited to single-layer networks Extends to multilayer networks
[70, 71] Multiplex. Fig. 1 Saddle-point dynamics: $\dot{y}=-\nabla_{y}\tilde{f}(y)-\mathcal{L}y-\mathcal{L}\lambda,$ $\dot{\lambda}=\mathcal{L}y$ Replica nodes required Removes need for replica nodes
[23, 26] Multilayer. Fig. 2 Tensor-based diffusion : $\frac{dX_{\beta\tilde{\nu}}(t)}{dt}=-\mathcal{L}^{\alpha\tilde{\mu}}_{\beta\tilde{\nu}}X_{\alpha\tilde{\mu}}(t)$ Focused on diffusion, not optimization Generalizes to distributed optimization
@@ II-A2 Multilayer Networks
are the fourth-order (i.e., rank-4) tensors of the canonical basis in the space $\mathbb{R}^{N\times N\times M\times M}$. We have adopted the notation $N\equiv N(M)$ to remark that the number of nodes in each layer might be different. The multilayer adjacency tensor $M^{\alpha\tilde{\mu}}_{\beta\tilde{\nu}}$ is a versatile construct that can represent a wide range of complex relationships between nodes across different layers.
@@ II-B Diffusion Processes on Multilayer Networks
II-B1 Diffusion Processes on Single-layer Networks
Using tensor notation, the diffusion processes on single-layer networks can be described as follows. Let $x_{\alpha}(t)$ be a state vector of nodes at time $t$. The diffusion equation is given by
@@ II-B1 Diffusion Processes on Single-layer Networks
II-B2 Diffusion Processes on Multilayer Networks
In this case, a diffusion equation for multilayer networks must incorporate terms that account for interlayer diffusion. Let $X_{\alpha\tilde{\mu}}(t)$ denote the state tensor of the nodes in each layer at time $t$. Thus, a diffusion equation for a multilayer network is obtained as follows,
@@ II-B2 Diffusion Processes on Multilayer Networks
Similar to the single-layer case, a multilayer combinatorial Laplacian tensor is presented as
@@ III Problem Statement
$f_{\alpha\tilde{\mu}}(x)=\sum_{\tilde{h}=1}^{M}\sum_{i=1}^{N}f_{i\tilde{h}}(x)e_{\alpha}(i)e_{\tilde{\mu}}(\tilde{h}),$ (22)
@@ VI Conclusion
In this paper, we develop a distributed optimization algorithm tailored for generalized multilayer networks within a tensorial framework. By extending the connection between diffusion and consensus dynamics and leveraging control approaches used in multiplex networks, we propose a distributed tensorial primal-dual saddle-point algorithm that achieves fast convergence and robustness. Similarly to multiplex networks, we find that intra- and interlayer diffusion constants serve as control parameters for optimal consensus time. Each layer first reaches a local consensus before achieving global consensus across layers.
@@
Possible extensions of this work involve obtaining a diffusion operator for more general higher-order networks and the application of the multilayer optimization proposed approach [73]. A natural continuation of this work is to study hypergraphs consensus dynamics [74] and consensus in simplicial complexes [75, 76]. On the other hand, topological Dirac operators offer a generalization scenario, which could be exploited to reveal new properties in complex systems, including new consensus dynamics [77]. Another potential extension involves the introduction of disorder [78, 79] and the application of graph zeta functions to analyze diffusion dynamics in higher-order networks [80].
@@ Appendix B Proof of Theorem 1
First, consider the convergence analysis for the tensor points $(Y^{\star}_{\beta\tilde{\nu}},\Lambda^{\star}_{\beta\tilde{\nu}})$. Let $\tilde{Y}_{\beta\tilde{\nu}}=Y_{\beta\tilde{\nu}}-Y^{\star}_{\beta\tilde{\nu}}$ and $\tilde{\Lambda}_{\beta\tilde{\nu}}=\Lambda_{\beta\tilde{\nu}}-\Lambda^{\star}_{\beta\tilde{\nu}}$. In the new variables, the saddle-point flow (26)-(27) yields
@@
To demonstrate the global asymptotic stability of the equilibrium tensor point $(Y^{\star},\Lambda^{\star})$, we apply LaSalle’s invariance principle, showing that there are no trajectories in $E$ other than the equilibrium points $\tilde{Y}=0$ and $\tilde{\Lambda}=0$. It is observed that if $\tilde{Y}=0$ the system (36)-(37) casts to
@@
Examining the dual dynamics (37), we can observe that $U^{\alpha\tilde{\mu}}\dot{\Lambda}_{\alpha\tilde{\mu}}=0$. Therefore, we obtain $\text{average}(\Lambda(t))=\text{average}(\Lambda_{0})$ $\forall t\geq 0$. As a consequence, the convergence to a single tensor saddle-point $(Y^{\star},\Lambda^{\star})$ of the flow (36)-(37) is guaranteed, which satisfies $Y^{\star}_{\alpha\tilde{\mu}}=x^{\star}U_{\alpha\tilde{\mu}}$ and $\Lambda^{\star}_{\alpha\tilde{\mu}}=\bar{\Lambda}_{\alpha\tilde{\mu}}+\text{average}(\Lambda_{0})U_{\alpha\tilde{\mu}}$. ∎